Journalist Report – January 30th

Journalist Report
Crew #: 328
Position: Crew Journalist, Comms Officer & Astronomer
Report prepared by: Tom Bickmore
Date: 30-01-2026
Sol: 5

Journalist Report Title: Sol 5

Hello from the Hab on Sol 5. A big day here at the Mars Desert Research Station: Our first live-delayed outreach session and a solid EVA, plus some nice progress inside.

Last night (Sol 4 evening) we had a movie night. It took surprisingly long to find a sci-fi film none of us had seen. We thought we had one, but halfway through two crew members realized they’d watched it years ago. Fun to see how the group reacted to the movie crew’s actions and reactions and to discuss what we’d handle differently in a similar scenario.

Today we ran our first Live-but-delayed outreach event: about a dozen kids from 4 countries joined us for a continuous conversation with the ~10-minute interplanetary comms delay simulated (constant video feeds both ways, delayed between parties). This was a rehearsal with our co-host students to test the full setup before the larger-audience version. We were all impressed by the kids’ questions, the depth of their knowledge, and how well they answered each other. Their variety of knowledge and their enthusiasm were incredible. The delay felt real and impactful: watching Earth respond to our "past" selves, then seeing our own delayed feed return. It made the interaction feel connected rather than isolated. Everyone had a great time.

We also had an EVA: Commander Mariló Torres and I went to the northern limit of our range, an area reported for signs of ancient life (it’s an active dinosaur quarry). No fossils or ancient traces found, but we did spot a very small live "dinosaur" (a lizard). Active dig sites had protective covers in place. The geology was striking. We saw distinct layers of rock with varied erosion patterns. On the drive to and from, we noticed crystalline patterns in the dirt where water gathers, looking a lot like the growths in Jahnavi’s soil experiments and similar to road-salt crystallization during Canadian thaws. Since it’s a desert dirt road with no added salt, these must be natural salts efflorescing from the soil. We plan to look more closely and sample these locations on a future EVA.

Commander Mariló Torres and HSO Aaron Tenner gave us a thorough safety briefing on additional Hab emergency procedures—good refresh and details.

A previously "delayed in orbit" supply drop arrived today with key chemicals and additives for Jahnavi’s soil experiments (to mix into local regolith for bricks and materials). This unblocks the main Science Dome work that was waiting on these items. Very exciting!

In the GreenHab, radish seeds are already germinating in the hydroponics setup. The other experiments (radish microgreens across four conditions, tomatoes in lab-regolith with 10% organic) continue progressing well.

Another day of progress with personal experiments planned for Sol 6!

EVA Report – January 30th

Report title: EVA Report
Crew #: Mariló Torres
Position: Commander
Report prepared by: Mariló Torres
Date: 30-01-2026
Sol: 5

EVA #: 4
Purpose of EVA:
Long-distance exploration with the objective of evaluating and surveying the area north of the station.
Start time: 14:02
End time: 16:13

Narrative:
Considering that Mission Support felt the Crew Scientist had not yet fully recovered, we decided to proceed with the EVA plan involving the Commander and the Crew Scientist. Therefore, we carried out the first EVA composed of only two crew members, the longest one to date, and the one that has involved the furthest distance traveled so far. And so it was. Aboard two rovers, we reached our objective and embarked on a trek through the canyon and the surrounding landscape.

In addition to analyzing the area and studying it for future EVAs, we documented the presence of reptiles and insects with photographs and found evidence of larger animals.

The positive aspects to highlight, once again, were the perfect execution of the operation, the coordination, and the punctuality. The negative aspect was that we saw several white vans on the way there and encountered several cars on the road on the way back, which somewhat disrupted our immersion in the simulation.

Destination per MDRS Map:
Along Cow Dung Road to Gateway to Lith and Special Region.

Destination Coordinates:
518000, 4256000

EVA Participants:
Commander and Crew Journalist.

Road(s) and routes per MDRS Map:
Cow Dung road 0110

Mode of travel:
By rover

GreenHab Report – January 30th

Report title: GreenHab Report
Crew #: 328
Position: Greenhab Officer
Report prepared by: Rebeca Goncalves
Date: 30/01/2026
Sol: 5

Environmental control (fan & heater): Fan max: 85F, Fan min: 80F, Heater: 70F
Average temperatures (last 24h): 75.9F
Maximum temperature (last 24h): 85F
Minimum temperature (last 24h): 66.8

Hours of supplemental light: 5:30h (6pm to 11:30pm)

Daily water usage for crops: 1
Daily water usage for research and/or other purposes: 0
Water in the Tank (160 gal useful capacity): 142 gallons remaining
Time(s) of watering for crops: 7pm

Changes to crops: seeds from hydroponics started to germinate

Narrative: watered crops and noted down germination rate for both hydroponics and soil system for Exp. 2. Tomatoes (Exp 1) have not yet germinated. Greenhouses remain covered.

Harvest: (include which crop and mass in grams) N/A

Support/supplies needed: 1 gallon of distilled water (from greenhab shelf). 0.5 gallons remain.

Sol Summary – January 30th

Report title: Sol Summary Report
Crew #: 328
Position: Commander
Report prepared by: Mariló Torres
Date: 30-01-2026
Sol: 5

Summary Title: Exploration
Mission Status: At full capacity and at full speed

Sol Activity Summary:
There is no more rewarding activity than inspiring new generations to love space. And that’s exactly what the online event that connected us with Earth aimed to do, answering questions from numerous children around the world. Thanks to our Crew Journalist, we were able to enjoy this heartwarming experience.

The Commander gave a safety briefing on the emergency plan in case of fire, and the HSO/Crew Engineer gave another on responding to a medical emergency.

After lunch, we embarked on our first EVA, consisting of only two crew members and the longest one to date. An opportunity to explore Mars and appreciate the magnificence of its landscape.

Look Ahead Plan:
Back to our experiments and scientific fieldwork. We’ll be spending many hours between the Green Hub, the Laboratory, and the RAM.

Anomalies in work:
None.

Weather:
Blue skies.

EVA:
Yes

Operations Report – January 30th

Report title: Operations Report
Crew #: 328
Position: Crew Engineer/HSO
Report prepared by: Aaron Tenner
Date: 30-01-2026
Sol: 5

Non-nominal systems: N/A
Notes on non-nominal systems: N/A
ROVERS
Spirit rover used: Yes
Hours: 292.2
Beginning charge: 100%
Ending charge: 41%
Currently charging: Yes
Opportunity rover used: Yes
Hours: 292.2
Beginning charge: 100%
Ending charge: 41%
Currently charging: Yes
Curiosity rover used: N/A
Hours:
Beginning charge:
Ending charge:
Currently charging:
Perseverance rover used: N/A
Hours:
Beginning charge:
Ending charge:
Currently charging:
General notes on rovers: Only spirit and opportunity were used today, both are now charging.
Summary of Hab operations:
Water Use (see notes): 25.45 gal/day
Main tank level (remaining gallons, see notes): 283.09 gal
Main water tank pipe heater (ON or OFF): ON
Main water tank heater (ON or OFF): ON
Toilet tank emptied (NO or YES): NO
Summary of internet: downlink 163.14 mbit/s, uplink 37.39 mbit/s
Summary of suits and radios: Suit 1 and 5 used today, replaced on chargers. Comms 1 and 5 used and returned after use.
Summary of GreenHab operations: (use information from the GreenHab report)
Water use: 2 gallons
Heater (ON or OFF): ON
Supplemental light (hours of operation): 6-11:30pm
Harvest (name, weight in grams): N/A
Summary of ScienceDome operations: Working on getting a microscope set up.
Dual split (Heat or AC, On or Off or Automatic): Automatic, trying to keep
temperature at 70-72 degrees F for regolith study.
Summary of RAM operations: None today.
Summary of any observatory issues: N/A
Summary of health and safety issues: Crew Engineer got a minor cut when putting up a flag.
Questions, concerns and requests to Mission Support:

Note: Crew wanted to express our deepest gratitude to mission support for all the support provided, thank you.

Journalist Report – January 29th

Journalist Report
Crew #: 328
Position: Crew Journalist, Comms Officer, & Astronomer
Report prepared by: Tom Bickmore
Date: 29-01-2026
Sol: 4

Journalist Report Title: Sol 4

Hello from the Hab on Sol 4. Another focused inside day at the Mars Desert Research Station—no EVAs, but steady progress on experiments and prep for tomorrow’s big activities.

I started early before sunrise, spotting the Chinese Space Station passing overhead (a nice real-world reminder of orbital activity), then captured a longer sunrise timelapse plus some pre-dawn shots with stars still visible—great light for photos.

We continued working on personal experiments throughout the day.

Crew Scientist Jahnavi Dangeti saw interesting developments in her soil testing: mineral crystals appear to be growing out of the brick samples she formed.
GreenHab Officer Rebeca Gonçalves finished testing her hydroponics setup and planted the radish seeds in it. The radishes are being grown as microgreens across four conditions for comparison: hydroponics, lab-certified regolith sim, locally collected regolith, and organic potting soil. Tomatoes are in the lab-regolith sim. All regolith-based setups include 10% organic potting soil by volume to support growth.
Engineer/Safety Officer Aaron Tenner worked on paperwork for his project, along with outreach tasks and planning.

I posted more of our daily reports and photos for outreach, got the Mars topography model (created and donated by Casey Handmer) running on my laptop after testing, and we’re awaiting confirmation to install it on the main Hab iMac. I also started setting up the Solar Telescope and preparing the Robotic Observatory for use—expanding my hands-on contributions beyond journaling.

Rebeca was assisted by Commander Mariló Torres on some of the GreenHab tending and planting work today.

This evening, with very clear skies, we watched the ISS (or our "Martian Space Station" equivalent) pass overhead—bright, steady, and a great view. I captured a Martian sunset timelapse to close out the day.

We’re gearing up for two key activities tomorrow: a longer EVA to the north, and our first Live-but-delayed session with kids on Earth. This one is a rehearsal with the same ~10-minute comms delay setup, constant video feeds (delayed between parties), and the same co-host students who will run the larger-audience event next time. It’s a chance to test the flow and tech before the full version.

Daily routines are settling in well, and the crew is staying productive.

More on how Sol 5 goes.

Operations Report – January 29th

Report title: Operations Report
Crew #: 328
Position: Crew Engineer/HSO
Report prepared by: Aaron Tenner
Date: (29-01-2026)
Sol: 4

Non-nominal systems: N/A
Notes on non-nominal systems: N/A
ROVERS
Spirit rover used: N/A
Hours:
Beginning charge:
Ending charge:
Currently charging:
Opportunity rover used: N/A
Hours:
Beginning charge:
Ending charge:
Currently charging:
Curiosity rover used: N/A
Hours:
Beginning charge:
Ending charge:
Currently charging:
Perseverance rover used: N/A
Hours:
Beginning charge:
Ending charge:
Currently charging:
General notes on rovers: Rovers were not used today.
Summary of Hab operations:
Water Use (see notes): 24.51 gal/day
Main tank level (remaining gallons, see notes): 308.54 gal
Main water tank pipe heater (ON or OFF): ON
Main water tank heater (ON or OFF): ON
Toilet tank emptied (NO or YES): NO
Summary of internet: Downlink 142.65 mbit/s, uplink 27.26 mbit/s, ping 20 ms
Summary of suits and radios: Not used today, sitting on chargers ready for use.
Summary of GreenHab operations: Crew Astrobiologist finished setting up her experiments today.
Water use: 3 gallons
Heater (ON or OFF): ON
Supplemental light (hours of operation): 6pm – 11:30pm
Harvest (name, weight in grams): N/A
Summary of ScienceDome operations: Took photos of regolith from yesterday and found that the distilled water caused salts and minerals to crystallize on the surface giving it a living look.
Dual split (Heat or AC, On or Off or Automatic): Automatic
Summary of RAM operations: None today.
Summary of any observatory issues: Opened observatory for the first time today and sent photos to mission support astronomers for approval of use.
Summary of health and safety issues: Previous ailments continue to improve.
Questions, concerns and requests to Mission Support: None today, have a good night mission support.

GreenHab Report – January 29th

Report title: GreenHab Report
Crew #: 328
Position: Greenhab Officer
Report prepared by: Rebeca Goncalves
Date: 29/01/2026
Sol: 4

Environmental control (fan & heater): Fan max: 85F, Fan min: 80F, Heater: 70F
Average temperatures (last 24h): 77.8F
Maximum temperature (last 24h): 84.4F
Minimum temperature (last 24h): 66.8F

Hours of supplemental light: 5:30h (6pm to 11:30pm)

Daily water usage for crops: 0
Daily water usage for research and/or other purposes: 3
Water in the Tank (160 gal useful capacity): 143 gallons remaining
Time(s) of watering for crops: 7pm

Changes to crops: N/A
Narrative: No germination recorded yet.
Harvest: (include which crop and mass in grams) N/A

Support/supplies needed: N/A

Photos attached.

Copyright © The Mars Society. All rights reserved. | Main Site